OpenStudy (anonymous):

find the domain and range for f(x)=sqr rt(1-x)

OpenStudy (anonymous):

The expression inside the radical can be greater or equal to zero. If the expression inside the radical is negative you get complex numbers. So to figure out what x can be set what is inside the radical greater than or equal to zero like this \[1 - x \ge 0\]Solve for x and you will get \[x \le 1\] so the domain is \[(-\infty,1]\] The range deals with the y values. Since this is a positive square root function the range is \[[0, \infty)\]